OpenStudy (anonymous):

y=x2+4.6x+4.7 y=3.3x-3.7 how do i solve?

OpenStudy (radar):

y=y then x2+4.6x+4.7=3.3x-3.7 \[x ^{2}+1.3x+8.4=0\]

OpenStudy (anonymous):

i got that far, i just dont know what to do to find the solutions?

OpenStudy (radar):

Multiply thru by 10 to eliminate decimals\[10x ^{2}+13x+84=0\]Use the quadratic formula. A=10, B=13, and C=84\[-13\pm \sqrt{169-4*10*84}\over2*10\]\[-13\pm \sqrt{-3191}\over20\]\[-13\pm56.49i \over 20\]\[-0.65\pm2.82i\]

OpenStudy (anonymous):

these are my answer choices: A. (–3.47, 3.48) and (2.17, –15.17) B. (–3.47, –15.17) and (–2.17, –10.88) C. (–3.47, –15.17) and (2.17, 3.48) D. no solution

OpenStudy (radar):

I would go with d. no solution as my solution is a complex results and maybe not applicable to your problem.